# Empowering Interdisciplinary Insights with Dynamic Graph Embedding Trajectories

### v1

@@ 2.1 Reddit Community Graphs Reveals Content Specialization, Content Diversity, and Echo Chambers
The trajectories of video-sharing subreddits (Fig. 1c) demonstrate diverse levels of specialization. Subreddits with a narrow focus on promoting YouTube videos and small channels, such as r/GetMoreViewsYT, r/YouTube_startups, r/AdvertiseYourVideos, r/SmallYoutubers, and r/YouTubeSubscribeBoost, move within a confined region, illustrating a high degree of content homogeneity within these subreddits as users simultaneously spread the same videos within multiple subreddits for better visibility. In contrast, general subreddits like r/videos display a greater diversity of content, as shown from their more expansive trajectories. These findings are supported by the numeric values of $\mathrm{Jaccard}_{100}$ (Supp. Fig.3), where the overlap between the nearest neighbors of each video-related subreddit in the embedding space in adjacent timestamps is high for video-promotion subreddits. Details for the metrics are in Supp.B.3.
@@
On the other hand, for sports-related communities (Fig. 1b), subreddits with specialized topics, such as r/nba (subreddit for the National Basketball Association), r/nfl (subreddit for the National Football League), r/MMA (subreddit for mixed martial arts), r/SquaredCircle (subreddit for professional wrestling), demonstrate similar levels of movements to more general subreddits like r/sports. Notably, r/nba has a large overlap with r/nfl, indicating that these two subreddits share similar audience, posts, and content sharing pattern. Both NBA and NFL feature team-based sports, high-profile athletes, strategies, and have regular seasons followed by playoff rounds that culminate in a championship event. In case of content sharing, many videos feature athletes or moments that have transcended their respective sports and gained widespread popularity, which is appreciated by fans of both basketball (NBA) and American football (NFL). Compared to subreddits focused on videos, sports-related subreddits display more variability among their neighboring subreddits in the embedding space, as evidenced by the $\operatorname{Jaccard}_{100}$ index values averaged on an annual basis (Supplementary Table 4).
@@ 2.3 Evolving Language and LGBTQ+ Acceptance: A Lexical Analysis of Societal Shifts
Additionally, Supp. Table 3 provides a comprehensive view of the top five words associated with each term in the embedding space over time. The words “happy” and “delighted” retain their consistent meanings across the years, serving as constants in the lexical landscape. However, the term “gay,” once widely employed to convey happiness before the 1960s, underwent a profound transformation when they were used to refer to homosexuality in the 1970s and acquired proximity with negative words such as “forlorn” and “ugly.” This lexical shift reflects the societal struggle to grapple with evolving perceptions of homosexuality. Furthermore, LGBT-related words, including “gay,” “homosexual,” and “lesbian,” exhibit strong associations with “clubs” and “dance” during the 1970s and 1980s. This phenomenon corresponds to the development of a distinctive LGBTQ+ culture and language during this era. Bars and dance clubs emerged as vital meeting places for the LGBTQ+ community, providing safe spaces for socialization, self-expression, and the formation of supportive networks [30]. It is crucial to acknowledge that the portrayal of LGBTQ+ characters and issues in popular culture largely perpetuated negative stereotypes and discriminatory portrayals during the examined period. This further entrenched negative attitudes within the general public, making societal acceptance and understanding a complex and arduous journey [31, 32].
@@
In Fig. 3b, the trajectories of individual countries reveal distinct patterns of economic development and stability. The United States, the United Kingdom, and Germany [^3] Germany has been listed in UN Comtrade as a single sovereign state since 1991, following German reunification in October 1990. demonstrate relatively stable and consistent trading status throughout the examined period (1988-2022). On the other hand, China’s trajectory moves between MAE and OAE, reflecting its prolonged period of economic development characterized by comprehensive domestic reforms, the lifting of price controls, and the liberalization of trade policies [36, 37]. Russia exhibits significant movements between OAE and EDE. Its trajectory predominantly shifted towards the EDE region during the period 1992-1998, coinciding with a substantial 40% contraction in GDP [3]. Starting from the early 2000s, Russia moves towards the region occupied by OAEs, including the four middle-sized developed countries Switzerland, Belgium, Sweden, and the Netherlands, which indicates a period of economic recovery characterized by greater trade volumes. Despite its status as an MAE, Japan has experienced economic development with significant fluctuations. The country encountered unique obstacles such as the Japanese asset price bubble (1990-1992) whose impact has lasted for more than a decade [38, 39]. We further use the Jaccard index [16] and Rank-biased Overlap (RBO) [17, 18] to measure the macro-level changes over time. Detailed calculations of these metrics are in Supplementary Section B.3. As reflected in Fig. 3c, the RBO and $\mathrm{Jaccard}_{5}$ for Japan plummeted during this period compared to other countries, indicating a period of instability in its economic status.
@@ 2.7 Modeling Social Dynamics in Ant Colonies on Animal Activity Graphs
Supplementary Fig. 5 illustrates these findings, showing that the movement patterns of nurses are generally restricted to areas near the queen, reflecting their frequent interactions. Conversely, foragers are typically found in remote areas, aligning with their external foraging activities and minimal contact with the queen. The spatial distribution and movements of these roles over time reveal distinct patterns: nurses and foragers maintain localized activity areas, whereas cleaners exhibit movement patterns intersecting with those of nurses due to their intermediary tasks.
@@ A Dataset Introduction
We used 9 publicly available datasets spanning 8 different domains to demonstrate DyGETViz’s wide applicability across all of these subject areas. Supp. Table 1 provides the statistics of the nine datasets.
@@
Supplementary Table 1: Statistics of our datasets. “Interval” indicates the time interval for each snapshot. ‘\’ indicates that the snapshot interval is not constant.
@@ B Method
Supplementary Figure 1: Our proposed DyGETViz framework.
In this section, we introduce our novel and computationally efficient framework DyGETViz for visualizing and analyzing dynamic graph embedding trajectories. Our framework effectively addresses the challenges associated with DGs mentioned in Section 1, including continuously evolving node embeddings and constant node addition and deletion. Figure 1 and Algorithm 1 describe the workflow and the pseudocode of DyGETViz, respectively.
@@ B.2 Embedding Visualization
Algorithm 1 DTDG embedding visualization. $\{\mathbf{V}^{t}\}$ is the set of temporal embedding matrix, where $\mathbf{V}^{t}\in\mathbb{R}^{|V^{t}|\times d}$. $\mathbf{X}\in\mathbb{R}^{|V^{\prime}|\times d}$ is the static embedding matrix for the anchor nodes. $\operatorname{sim}(\cdot):\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$ is a similarity measure. $\mathcal{N}(i,k,t)$ denotes the $k$ nearest neighbors of $v_{i}$ at time $t$ in the embedding space. $\operatorname{Agg}(\cdot)$ is an aggregation function. $\alpha$ is an interpolation factor. $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{|V^{\prime}|}\in\mathbb{R}^{|V^{\prime}|\times p}$ is the $p$-dimensional projection of nodes $v_{i}\in V^{\prime}$.
@@
It is worth noting that alternative ranking evaluation measures, such as Spearman’s rank correlation coefficient [52] and Kendall’s tau [53, 54], exist. However, these measures do not explicitly differentiate the importance of the ranks at different positions in the list and are sensitive to small perturbations of rankings, particularly towards the middle of the list [55]. To demonstrate this, Supp. Table 4 shows the distribution of average cosine similarity for all nodes in the four datasets, HistWords-CN, Reddit, Ant, and DGraph. We observe that the cosine similarity usually plateaus in the middle range, suggesting a large number of nodes with highly similar cosine similarity.
@@ C.2 Visualization
Supplementary Figure 2: Comparison of varying $f(\cdot)$ (Line 6) and $k$ (Line 12) in Algorithm 1.
Supplementary Table 2: Notations used in this paper
@@
Supplementary Figure 3: $\operatorname{Jaccard}_{100}$ among the subreddits in terms of videos shared between them.
Supplementary Figure 4: Distribution of average cosine similarity of the four datasets: HistWords-CN, Reddit, Ant, and DGraph. The x- and y- axis represents the percentage of the nodes in the dataset and cosine similarities, respectively.
Supplementary Table 3: Evolution of word associations from the 1950s to the 1990s. The term “gay” initially carried connotations with happiness and fortune, but underwent a decline in positivity during the 1970s as its association with homosexuality became more prevalent.
@@
Supplementary Table 4: Average $\operatorname{Jaccard}_{100}$ of subreddits in each year.
@@ D Additional Experimental Results
Supplementary Figure 5: Trajectories of ants from the Ant dataset [9].
@@ D.1 Enron: Email Communication
Supplementary Figure 6: Visualization of the Enron dataset [44]. (a) RBO and $\operatorname{Jaccard}_{3}$ during the projection time period. (b) Dynamic visualization of the node trajectories. The positions of employees are annotated. Employees occupying managerial/president positions (Mike Grigsby, Louise Kitchen and Greg Whalley) other than the CEOs interact with employees at various levels, resulting in more diverse trajectories compared to CEOs (Kenneth Lay and Jeffrey Skilling) and regular employees (e.g.Scott Neal). Furthermore, we observe a decline in both RBO and $\operatorname{Jaccard}_{3}$ from January to March 2001, suggesting a significant shift in communication partners among all employees. This decline aligns with the CEO transition in February 2001.
@@
In Fig. 6b, the trajectories of Lay and Skilling indicate relatively static communication communities, confined within a small range that mainly involves a limited number of vice presidents and managers. These patterns reflect their shortcomings in two-way communication, as demonstrated in previous studies [66, 67] — the failure to deliver honest, ethical messages to employees and the lack of awareness regarding company operations. Meanwhile, ordinary employees such as Geir Solberg, Kay Mann, and Scott Neal have specific job functions that confine their communication to their respective teams, resulting in relatively fixed communication patterns with a limited number of partners and trajectories with less variability.
@@
Moreover, as shown in Fig. 6a, the decline in RBO and $\operatorname{Jaccard}_{3}$ across all employees during the CEO transition from Lay to Skilling (January - March 2001) highlights the impact of leadership changes on communication dynamics within the organization. DyGETViz uncovers differences in communication patterns among employees, offering a novel perspective on organizational structure and dynamics. Understanding these patterns provides valuable insights for future research and organizational management.
@@ D.2 Dynamic Temporal and Spatial Modeling of Chickenpox Spread in Hungary
Supplementary Figure 7: a. Trajectories of the 19 counties and the capital Budapest in the Chickenpox dataset [45]. The most (least) populous counties are plotted using colors to the red (purple) side of the spectrum, and vice versa. b. Embedding movements of the 20 nodes. c/d. Population and average weekly chickenpox cases of all counties.
@@
In this study, we use the Hungary Chickenpox dataset [45], which includes the weekly chickenpox cases in Hungarian counties and the capital Budapest between 2005 and 2015. From the trajectories in Fig. 7a, we found that the capital city Budapest stands out as the node with the most movements due to its high population and . As the second most populous county in the country, Pest exhibits a trajectory that significantly overlaps with Budapest at each snapshot, indicating a high seasonality in the number of cases. This overlap can be attributed to factors such as geographical locations and suburbanization in the metropolitan area of Budapest, which caused considerable population movements between the two regions. According to a census in 2011 [71], nearly 60% of commuters living in the suburban zone of Pest work in Budapest. Such population overlap can facilitate the spread of diseases. Bács-Kiskun, a bordering state of Pest, moves towards Pest in winter, especially the middle of December, but away from it in summer, indicating a periodicity between the winter surge and the summer decline [45, 72]. In contrast, counties such as Tolna, Vas, Zala, and Heves, which are among the five least populated counties, form their own clusters with movements limited to the lower left of the plot, indicating a lower susceptibility to diseases due to smaller populations and fewer demographic movements.

### v2

@@ 2.1 Reddit Community Graphs Reveals Content Specialization, Content Diversity, and Echo Chambers
The trajectories of video-sharing subreddits (Fig. 1c) demonstrate diverse levels of specialization. Subreddits with a narrow focus on promoting YouTube videos and small channels, such as r/GetMoreViewsYT, r/YouTube_startups, r/AdvertiseYourVideos, r/SmallYoutubers, and r/YouTubeSubscribeBoost, move within a confined region, illustrating a high degree of content homogeneity within these subreddits as users simultaneously spread the same videos within multiple subreddits for better visibility. In contrast, general subreddits like r/videos display a greater diversity of content, as shown from their more expansive trajectories. These findings are supported by the numeric values of $\mathrm{Jaccard}_{100}$ (Fig. S3), where the overlap between the nearest neighbors of each video-related subreddit in the embedding space in adjacent timestamps is high for video-promotion subreddits. Details for the metrics are in Supplementary Sec. B.3.
@@
On the other hand, for sports-related communities (Fig. 1b), subreddits with specialized topics, such as r/nba (subreddit for the National Basketball Association), r/nfl (subreddit for the National Football League), r/MMA (subreddit for mixed martial arts), r/SquaredCircle (subreddit for professional wrestling), demonstrate similar levels of movements to more general subreddits like r/sports. Notably, r/nba has a large overlap with r/nfl, indicating that these two subreddits share similar audience, posts, and content sharing pattern. Both NBA and NFL feature team-based sports, high-profile athletes, strategies, and have regular seasons followed by playoff rounds that culminate in a championship event. In case of content sharing, many videos feature athletes or moments that have transcended their respective sports and gained widespread popularity, which is appreciated by fans of both basketball (NBA) and American football (NFL). Compared to subreddits focused on videos, sports-related subreddits display more variability among their neighboring subreddits in the embedding space, as evidenced by the $\operatorname{Jaccard}_{100}$ index values averaged on an annual basis (Table S4).
@@ 2.3 Evolving Language and LGBTQ+ Acceptance: A Lexical Analysis of Societal Shifts
Additionally, Table S3 provides a comprehensive view of the top five words associated with each term in the embedding space over time. The words “happy” and “delighted” retain their consistent meanings across the years, serving as constants in the lexical landscape. However, the term “gay,” once widely employed to convey happiness before the 1960s, underwent a profound transformation when they were used to refer to homosexuality in the 1970s and acquired proximity with negative words such as “forlorn” and “ugly.” This lexical shift reflects the societal struggle to grapple with evolving perceptions of homosexuality. Furthermore, LGBT-related words, including “gay,” “homosexual,” and “lesbian,” exhibit strong associations with “clubs” and “dance” during the 1970s and 1980s. This phenomenon corresponds to the development of a distinctive LGBTQ+ culture and language during this era. Bars and dance clubs emerged as vital meeting places for the LGBTQ+ community, providing safe spaces for socialization, self-expression, and the formation of supportive networks [30]. It is crucial to acknowledge that the portrayal of LGBTQ+ characters and issues in popular culture largely perpetuated negative stereotypes and discriminatory portrayals during the examined period. This further entrenched negative attitudes within the general public, making societal acceptance and understanding a complex and arduous journey [31, 32].
@@
In Fig. 3b, the trajectories of individual countries reveal distinct patterns of economic development and stability. The United States, the United Kingdom, and Germany [^3] Germany has been listed in UN Comtrade as a single sovereign state since 1991, following German reunification in October 1990. demonstrate relatively stable and consistent trading status throughout the examined period (1988-2022). On the other hand, China’s trajectory moves between MAE and OAE, reflecting its prolonged period of economic development characterized by comprehensive domestic reforms, the lifting of price controls, and the liberalization of trade policies [36, 37]. Russia exhibits significant movements between OAE and EDE. Its trajectory predominantly shifted towards the EDE region during the period 1992-1998, coinciding with a substantial 40% contraction in GDP [3]. Starting from the early 2000s, Russia moves towards the region occupied by OAEs, including the four middle-sized developed countries Switzerland, Belgium, Sweden, and the Netherlands, which indicates a period of economic recovery characterized by greater trade volumes. Despite its status as an MAE, Japan has experienced economic development with significant fluctuations. The country encountered unique obstacles such as the Japanese asset price bubble (1990-1992) whose impact has lasted for more than a decade [38, 39]. We further use the Jaccard index [16] and Rank-biased Overlap (RBO) [17, 18] to measure the macro-level changes over time. Detailed calculations of these metrics are in Supplementary Sec. B.3. As reflected in Fig. 3c, the RBO and $\mathrm{Jaccard}_{5}$ for Japan plummeted during this period compared to other countries, indicating a period of instability in its economic status.
@@ 2.7 Modeling Social Dynamics in Ant Colonies on Animal Activity Graphs
Fig. S5 illustrates these findings, showing that the movement patterns of nurses are generally restricted to areas near the queen, reflecting their frequent interactions. Conversely, foragers are typically found in remote areas, aligning with their external foraging activities and minimal contact with the queen. The spatial distribution and movements of these roles over time reveal distinct patterns: nurses and foragers maintain localized activity areas, whereas cleaners exhibit movement patterns intersecting with those of nurses due to their intermediary tasks.
@@ 3 Conclusion and Future Works
Acknowledgment
This research/material is based upon work supported in part by NSF grants CNS-2154118, IIS-2027689, ITE-2137724, ITE-2230692, CNS2239879, Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112290102 (subcontract No. PO70745), and funding from Microsoft, Google, and Adobe Inc. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the position or policy of DARPA, DoD, SRI International, NSF and no official endorsement should be inferred. We thank the reviewers for their comments.
@@ A Dataset Introduction
We used 9 publicly available datasets spanning 8 different domains to demonstrate DyGETViz’s wide applicability across all of these subject areas. Table S1 provides the statistics of the nine datasets.
@@
Table S1: Statistics of our datasets. “Interval” indicates the time interval for each snapshot. ‘\’ indicates that the snapshot interval is not constant.
@@ B Method
Figure S1: Our proposed DyGETViz framework.
In this section, we introduce our novel and computationally efficient framework DyGETViz for visualizing and analyzing dynamic graph embedding trajectories. Our framework effectively addresses the challenges associated with DGs mentioned in Section 1, including continuously evolving node embeddings and constant node addition and deletion. Figure S1 and Algorithm S1 describe the workflow and the pseudocode of DyGETViz, respectively.
@@ B.2 Embedding Visualization
Algorithm S1 DTDG embedding visualization. $\{\mathbf{V}^{t}\}$ is the set of temporal embedding matrix, where $\mathbf{V}^{t}\in\mathbb{R}^{|V^{t}|\times d}$. $\mathbf{X}\in\mathbb{R}^{|V^{\prime}|\times d}$ is the static embedding matrix for the anchor nodes. $\operatorname{sim}(\cdot):\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$ is a similarity measure. $\mathcal{N}(i,k,t)$ denotes the $k$ nearest neighbors of $v_{i}$ at time $t$ in the embedding space. $\operatorname{Agg}(\cdot)$ is an aggregation function. $\alpha$ is an interpolation factor. $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{|V^{\prime}|}\in\mathbb{R}^{|V^{\prime}|\times p}$ is the $p$-dimensional projection of nodes $v_{i}\in V^{\prime}$.
@@
It is worth noting that alternative ranking evaluation measures, such as Spearman’s rank correlation coefficient [52] and Kendall’s tau [53, 54], exist. However, these measures do not explicitly differentiate the importance of the ranks at different positions in the list and are sensitive to small perturbations of rankings, particularly towards the middle of the list [55]. To demonstrate this, Supp. Table S4 shows the distribution of average cosine similarity for all nodes in the four datasets, HistWords-CN, Reddit, Ant, and DGraph. We observe that the cosine similarity usually plateaus in the middle range, suggesting a large number of nodes with highly similar cosine similarity.
@@ C.2 Visualization
Figure S2: Comparison of varying $f(\cdot)$ (Line 6) and $k$ (Line 12) in Algorithm S1.
Table S2: Notations used in this paper
@@
Figure S3: $\operatorname{Jaccard}_{100}$ among the subreddits in terms of videos shared between them.
Figure S4: Distribution of average cosine similarity of the four datasets: HistWords-CN, Reddit, Ant, and DGraph. The x- and y- axis represents the percentage of the nodes in the dataset and cosine similarities, respectively.
Table S3: Evolution of word associations from the 1950s to the 1990s. The term “gay” initially carried connotations with happiness and fortune, but underwent a decline in positivity during the 1970s as its association with homosexuality became more prevalent.
@@
Table S4: Average $\operatorname{Jaccard}_{100}$ of subreddits in each year.
@@ D Additional Experimental Results
Figure S5: Trajectories of ants from the Ant dataset [9].
@@ D.1 Enron: Email Communication
Figure S6: Visualization of the Enron dataset [44]. (a) RBO and $\operatorname{Jaccard}_{3}$ during the projection time period. (b) Dynamic visualization of the node trajectories. The positions of employees are annotated. Employees occupying managerial/president positions (Mike Grigsby, Louise Kitchen and Greg Whalley) other than the CEOs interact with employees at various levels, resulting in more diverse trajectories compared to CEOs (Kenneth Lay and Jeffrey Skilling) and regular employees (e.g.Scott Neal). Furthermore, we observe a decline in both RBO and $\operatorname{Jaccard}_{3}$ from January to March 2001, suggesting a significant shift in communication partners among all employees. This decline aligns with the CEO transition in February 2001.
@@
In Fig. S6b, the trajectories of Lay and Skilling indicate relatively static communication communities, confined within a small range that mainly involves a limited number of vice presidents and managers. These patterns reflect their shortcomings in two-way communication, as demonstrated in previous studies [66, 67] — the failure to deliver honest, ethical messages to employees and the lack of awareness regarding company operations. Meanwhile, ordinary employees such as Geir Solberg, Kay Mann, and Scott Neal have specific job functions that confine their communication to their respective teams, resulting in relatively fixed communication patterns with a limited number of partners and trajectories with less variability.
@@
Moreover, as shown in Fig. S6a, the decline in RBO and $\operatorname{Jaccard}_{3}$ across all employees during the CEO transition from Lay to Skilling (January - March 2001) highlights the impact of leadership changes on communication dynamics within the organization. DyGETViz uncovers differences in communication patterns among employees, offering a novel perspective on organizational structure and dynamics. Understanding these patterns provides valuable insights for future research and organizational management.
@@ D.2 Dynamic Temporal and Spatial Modeling of Chickenpox Spread in Hungary
Figure S7: a. Trajectories of the 19 counties and the capital Budapest in the Chickenpox dataset [45]. The most (least) populous counties are plotted using colors to the red (purple) side of the spectrum, and vice versa. b. Embedding movements of the 20 nodes. c/d. Population and average weekly chickenpox cases of all counties.
@@
In this study, we use the Hungary Chickenpox dataset [45], which includes the weekly chickenpox cases in Hungarian counties and the capital Budapest between 2005 and 2015. From the trajectories in Fig. S7a, we found that the capital city Budapest stands out as the node with the most movements due to its high population and . As the second most populous county in the country, Pest exhibits a trajectory that significantly overlaps with Budapest at each snapshot, indicating a high seasonality in the number of cases. This overlap can be attributed to factors such as geographical locations and suburbanization in the metropolitan area of Budapest, which caused considerable population movements between the two regions. According to a census in 2011 [71], nearly 60% of commuters living in the suburban zone of Pest work in Budapest. Such population overlap can facilitate the spread of diseases. Bács-Kiskun, a bordering state of Pest, moves towards Pest in winter, especially the middle of December, but away from it in summer, indicating a periodicity between the winter surge and the summer decline [45, 72]. In contrast, counties such as Tolna, Vas, Zala, and Heves, which are among the five least populated counties, form their own clusters with movements limited to the lower left of the plot, indicating a lower susceptibility to diseases due to smaller populations and fewer demographic movements.